# Permutation-Invariant Spectral Learning via Dyson Diffusion

### v1

@@ Appendix B Spectral Dyson SDE
Theorem 3.1 (restated).
Denote by $\lambda(t)=(\lambda_{1}(t),\ldots,\lambda_{n}(t))$ the ordered spectrum of the matrix-valued Ornstein-Uhlenbeck process $M(t)$ of eq. 1. Then assuming that the initial matrix $M(0)$ has simple spectrum, $\lambda(t)$ satisfies for all $1\leq k\leq n$ the stochastic differential equation
@@
for $W_{1},\ldots,W_{n}$ independent standard Brownian motions. Moreover the unique stationary distribution of (Dyson-BM) has density
$p_{\mathrm{inv}}(\lambda)=\frac{1}{Z}\exp(-U(\lambda))\quad\quad\text{ for }\quad\quad U(\lambda)=\frac{\beta}{2\alpha}\sum_{k}\lambda_{k}^{2}-\sum_{k<\ell}\ln|\lambda_{k}-\lambda_{\ell}|,$
for $\lambda\in C_{n}$ and $Z$ a normalizing constant so that $p_{\mathrm{inv}}$ corresponds to a probability measure.
@@
We first prove Dyson-BM. This part of the proof is based on Keating (2023) but generalizes it to arbitrary coefficients. For mathematical details on the $\alpha=\frac{1}{n}$, $\beta=0$ case, see also Anderson et al. (2009). Suppose $M(t)$ satisfies the SDE eq. 1. That is, with $M(0)\in\mathrm{Sym}(\mathbb{R}^{n\times n})$ we have
@@
where the first part is 0 since $\lambda_{k}(t)$ is only a function of the $M_{\eta}(t)$, not of time. By eq. 1, we get
@@
Multiplying eq. 22 by $v_{l}^{T}$ from the left yields
@@
$(I)$ $=(1-\delta_{kl})\lambda_{l}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}$ (by eq. 25 )
@@
$(III)$ $=(1-\delta_{kl})\lambda_{k}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}},$ (by eq. 25 )
@@
Using that the $v_{k}$ are an ONB of eigenvectors, we can conclude with eq. 32 that the first summand of eq. 20 is
@@
which yields with eq. 33
@@
$\frac{\partial^{2}\lambda_{k}}{\partial M_{ij}\partial M_{ij}}=$ $\frac{\partial}{\partial M_{ij}}\left((v_{k})_{i}(v_{k})_{j}(2-\delta_{ij})\right)$ (by eq. 32)
@@
These second order partial derivative are summed over $\mathcal{I}$ in eq. 20, so that combining eq. 38 with the definition of $D_{ij}$ yields
@@
where we realize that upon accounting for all $\delta_{ij}$, the first two summands are simply $\norm{v_{l}}_{2}^{2}\norm{v_{k}}_{2}^{2}$, while the third summand may we written an inner product
@@
It remains to determine an explicit expression of the Brownian motion in eq. 20. We have
@@
We can thus conclude that by eq. 20 we have,
@@
for $\lambda\in C_{n}$ and $Z$ a normalizing constant so that $p_{\mathrm{inv}}$ is a probability measure, is the invariant distribution. We use the Fokker-Plank equation. Indeed, recall that if we have an SDE
@@
Figure 4: Plot of the invariant density of Dyson-BM for $d=2$ and $\alpha=\beta=1$
@@ Appendix C Inferring the Eigenvector Dynamics
Theorem 3.2 (restated).
Denote by $(v_{1}(t),\ldots,v_{n}(t))$ the orthonormal eigenvectors associated to the eigenvalues of Theorem 3.1. Assuming that the initial matrix $M(0)$ has simple spectrum, $v_{k}(t)$ satisfies for $k\in[n]$ the stochastic differential equation
@@
Analogously to the proof of Theorem 3.1, we may view the $v_{k}$ for $k\in[n]$ as a function of the matrix components $M_{ij}$ for $(i,j)\in\mathcal{I}\coloneqq\{(i,j):1\leq i\leq j\leq n\}$. We thus have by Itô’s lemma
@@
For the first summand of eq. 48, we observe, using eq. 36, that
@@
For the second summand of eq. 48, we observe
@@
and using Equation 30 and Equation 36 we get
@@
For term (ii) we have by Equation 30 and Equation 36
@@
By using Equation 30 and Equation 36, we get
@@
By pulling in the $\eta$-dependent terms from Equation 48 which affect term $(iv)$, and then by subsequently pulling the sum over $\eta$ in, we get
@@
which, using eq. 58, becomes
@@
Since the linear combination of independent Brownian Motions is jointly normal, we see from the Covariance property in eq. 60 that $\tilde{B}_{lk}$ is independent of $\tilde{B}_{ab}$ for $(l,k)\neq(a,b)$ and $(l,k)\neq(b,a)$. ∎
We can now turn to the third summand of eq. 48.
@@
where we used Lemma C.1 in the last step so that $\mathrm{d}\tilde{B}_{lk}$ are symmetric Brownian motions with variance $1+\delta_{lk}$.
Having established all the terms in the SDE in Theorem 3.2, we check that this dynamics of the eigenvectors gives rise to normalized vectors, assuming that the initial vectors $(v_{1}(0),\dots,v_{n}(0))$ are normalized. To this end, it suffices to note that $v_{k}(t+\mathrm{d}t)=v_{k}(t)+dv_{k}(t)$ is normalized given that $v_{k}(t)$ is normalized, that is $(v_{k}(t))^{2}=1$ for any $t$. By continuity of $t\mapsto v_{k}(t)$, it suffices to show that $v_{k}(\mathrm{d}t)=v_{k}(0)+dv_{k}(0)$ remains normalized. We use the SDE from Theorem 3.2 to compute that up terms of the order of $\mathrm{d}t^{3/2}$ or higher we have (omitting “(0)” in the notation)
@@
The argument may be iterated and thus the claim hold for all $t>0$.
@@ Appendix D Time reversal: Existence and Uniqueness
The time-reversal of the Dyson-BM in the sense of Ref. Anderson (1982) is given in Eq. (5). Since the drift coefficient in the Dyson-BM is not locally Lipschitz continuous, the existence and uniqueness of strong solutions to Dyson-BM and Eq. (5), and the applicability of Ref. Anderson (1982) are not obvious. While the existence and uniqueness of a strong solution to the Dyson-BM is well established (see, e.g., Lemma 4.3.3 in Ref. Anderson et al. (2009)), we here sketch how to ensure the other two points.
To ensure existence and uniqueness of a strong solution to Eq. (5), repeat the arguments in Lemma 4.3.3 in Ref. Anderson et al. (2009), where the divergent $x^{-1}$-term in the drift is replaced by the locally Lipschitz continuous approximation $\phi(x)=x^{-1}$ for $|x|\geq R^{-1}$ and $\phi(x)=xR^{2}$ otherwise. For any $R>0$, the desired statements follow from the local Lipschitz continuity of the drift, and details on the limit $R\to 0$ can be found in Ref. Anderson et al. (2009). The only difference to the forward motion is the additional drift term containing the score function. However, since the dynamics is almost surely contained in the interior of the Weyl chamber (Katori & Tanemura, 2003), the propagator in the score-contribution is, as usual, dominated by white noise as $\mathrm{d}t\to 0$. Therefore, this term will not cause complications as $R\to 0$ and the arguments from Ref. Anderson et al. (2009) imply uniqueness and existence of a strong solution to Eq. (5).
Using the same regularization of the divergent term, for any $R>0$ the statements of Ref. Anderson (1982) are directly applicable. Since this regularization is piecewise continuous, we can take the limit $R\to 0$ under time reversal. Since we just established the existence of the solution to Eq. (5), the limit $R\to 0$ converges. Furthermore, the arguments are expected to generalize (regularize $\phi(x)=x^{-2}$) to the eigenvector equations and their time reversal.
@@ Appendix E Time rescaling
We rescale time for Dyson-BM. Let $T(s)$ be a continuous, differentiable rescaling of time, monotonically increasing, with $T(0)=0$ (for convenience).
Let $\gamma_{k}(t):=\lambda(T(t))$. With this definition, the Ito SDE eq. Dyson-BM corresponds to the Ito integral
@@
$=$ $\int_{0}^{t}\left(\alpha\sum_{i\neq k}\frac{1}{\lambda_{k}(T(s))-\lambda_{i}(T(s))}-\beta\lambda_{k}(T(s))\right)T^{\prime}(s)\mathrm{d}s+\int_{0}^{t}\sqrt{2\alpha T^{\prime}(s)}\mathrm{d}B_{k}(s)$ (using Thm. 8.5.7 in Øksendal (2003))
@@
Dyson’s conjecture says that the $\lambda(\frac{1}{N},\frac{1}{2})$-SDE converges to global equilibrium in time $\Theta(1)$ (see Yang (2022)). Running $\lambda(\frac{1}{N},\frac{1}{2})$ until time 1 is the same as running the $\gamma(\frac{N}{2})$-SDE until time $T(1)=\frac{1}{N}$.
@@ Appendix F Stepsize Controller
Dyson Brownian Motion almost surely never crosses the singularities. Hence, conditioning on non-crossing corresponds to conditioning on a probability $1$ event, which does not change the dynamics. Given the noise, we can thus calculate the maximal step size, beyond which we would cross the singularity. This is a very useful upper bound, which we employ in practice to get the numerical scheme working. It has two effects: (1) close to the boundary of the Weyl Chamber, it avoids numerically stepping over the singularities and (2) far from the boundary, it allows for larger step size, increasing efficiency.
@@ F.1 Forward in Time
3:   $\delta t_{k}\leftarrow$ maximal step size based on $\lambda_{k+1}-\lambda_{k}$ and samples $u_{k}$, $u_{k+1}$ as described in Section F.1
@@
Figure 5: Forward step size controller which exploits that non-crossing of paths happens with probability 11. The exact calculations are carried out in Section F.1.
@@ F.2 Backward in Time
Backwards in time, we carry out the analogous computation for the more involved backwards SDE in eq. 5. In essence, this boils again down to solving a quadratic equation and considering all edge cases.
@@ Appendix G Shooting mechanism
In the forward dynamics, if a certain step size would lead to the probability-0 event of leaving the Weyl-Chamber, we know that the source of the error is the finite-time-step discrete approximation, so that decreasing the step size will always provide a fix. In the backward dynamics, however, this is not guaranteed: As described in Section 3.2, a possible reason for this probability-0 event in the backwards dynamics is that the score $s_{\theta}$ might be not perfectly learned: $s_{\theta}(\lambda,t)\neq s(\lambda,t)$ for some $\lambda,t$. To avoid this probability-0 event, we use the following “shooting mechanism”: If we were to leave the Weyl-Chamber, we replace the learned score with the analytically known score in the invariant state, eliminating the use of the neural network at that point and leading effectively to a repulsion with the negative forward drift. This mechanism is not expected to change the dynamics in any unfavorable way since it is only applied very rarely, and only in cases where the actual learned dynamics is a much worse approximation (since it would give rise to measure zero events).
Appendix H Inference

### v2

@@ Appendix B Spectral Dyson SDE
Theorem 2 (restated).
Denote by $\lambda(t)=(\lambda_{1}(t),\ldots,\lambda_{n}(t))$ the ordered spectrum of the matrix-valued Ornstein-Uhlenbeck process $M(t)$ of Equation˜1. Then assuming that the initial matrix $M(0)$ has simple spectrum, $\lambda(t)$ satisfies for all $1\leq k\leq n$ the stochastic differential equation
@@
for $W_{1},\ldots,W_{n}$ independent standard Brownian motions. Moreover the unique stationary distribution of (Dyson-BM) has density $p_{\mathrm{inv}}(\lambda)=\exp(-U(\lambda))/Z$ for $U(\lambda)=\frac{\beta}{2\alpha}\sum_{k}\lambda_{k}^{2}-\sum_{k<\ell}\ln|\lambda_{k}-\lambda_{\ell}|$, $\lambda\in C_{n}$, and $Z$ a normalizing constant so that $p_{\mathrm{inv}}$ corresponds to a probability measure.
@@
We first prove (Dyson-BM). This part of the proof is based on [46] but generalizes it to arbitrary coefficients. For mathematical details on the $\alpha=\frac{1}{n}$, $\beta=0$ case, see also [47]. Suppose $M(t)$ satisfies the SDE Equation˜1. That is, with $M(0)\in\mathrm{Sym}(\mathbb{R}^{n\times n})$ we have
@@
where the first part is 0 since $\lambda_{k}(t)$ is only a function of the $M_{\eta}(t)$, not of time. By Equation˜1, we get
@@
Multiplying Equation˜21 by $v_{l}^{T}$ from the left yields
@@
$(I)$ $=(1-\delta_{kl})\lambda_{l}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}$ (by Equation 24 )
@@
$(III)$ $=(1-\delta_{kl})\lambda_{k}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}},$ (by Equation 24 )
@@
Using that the $v_{k}$ are an ONB of eigenvectors, we can conclude with Equation˜31 that the first summand of Equation˜19 is
@@
which yields with Equation˜32
@@
$\frac{\partial^{2}\lambda_{k}}{\partial M_{ij}\partial M_{ij}}=$ $\frac{\partial}{\partial M_{ij}}\left((v_{k})_{i}(v_{k})_{j}(2-\delta_{ij})\right)$ (by Equation 31)
@@
These second order partial derivative are summed over $\mathcal{I}$ in Equation˜19, so that combining Equation˜37 with the definition of $D_{ij}$ yields
@@
where we realize that upon accounting for all $\delta_{ij}$, the first two summands are simply $\norm{v_{l}}_{2}^{2}\norm{v_{k}}_{2}^{2}$, while the third summand may be written an inner product
@@
It remains to determine an explicit expression of the Brownian motion in Equation˜19. We have
@@
We can thus conclude that by Equation˜19 we have,
@@
for $\lambda\in C_{n}$ and $Z$ a normalizing constant so that $p_{\mathrm{inv}}$ is a probability measure, is the invariant distribution. We use the Fokker-Planck equation. Indeed, recall that if we have an SDE
@@
Figure 4: Plot of the invariant density of (Dyson-BM) for $d=2$ and $\alpha=\beta=1$
@@ Appendix C Inferring the Eigenvector Dynamics
Theorem 3 (restated).
Denote by $(v_{1}(t),\ldots,v_{n}(t))$ a tuple of orthonormal eigenvectors associated to the eigenvalues of Theorem˜2. Assuming that the initial matrix $M(0)$ has simple spectrum, $v_{k}(t)$ satisfies for $k\in[n]$ the stochastic differential equation
@@
Analogously to the proof of Theorem˜2, we may view the $v_{k}$ for $k\in[n]$ as a function of the matrix components $M_{ij}$ for $(i,j)\in\mathcal{I}\coloneqq\{(i,j):1\leq i\leq j\leq n\}$. We thus have by Itô’s lemma
@@
For the first summand of Equation˜47, we observe, using Equation˜35, that
@@
For the second summand of Equation˜47, we observe
@@
and using Equation˜29 and Equation˜35 we get
@@
For term (ii) we have by Equation˜29 and Equation˜35
@@
By using Equation˜29 and Equation˜35, we get
@@
By pulling in the $\eta$-dependent terms from Equation˜47 which affect term $(iv)$, and then by subsequently pulling the sum over $\eta$ in, we get
@@
which, using Equation˜57, becomes
@@
Since the linear combination of independent Brownian motions is jointly normal, we see from the Covariance property in Equation˜59 that $\tilde{B}_{lk}$ is independent of $\tilde{B}_{ab}$ for $(l,k)\neq(a,b)$ and $(l,k)\neq(b,a)$. ∎
We can now turn to the third summand of Equation˜47.
@@
where we used Lemma˜5 in the last step so that $\mathrm{d}\tilde{B}_{lk}$ are symmetric Brownian motions with variance $1+\delta_{lk}$.
Having established all the terms in the SDE in Theorem 3, we check that this dynamics of the eigenvectors gives rise to normalized vectors, assuming that the initial vectors $(v_{1}(0),\dots,v_{n}(0))$ are normalized. To this end, it suffices to note that $v_{k}(t+\mathrm{d}t)=v_{k}(t)+dv_{k}(t)$ is normalized given that $v_{k}(t)$ is normalized, that is $(v_{k}(t))^{2}=1$ for any $t$. By continuity of $t\mapsto v_{k}(t)$, it suffices to show that $v_{k}(\mathrm{d}t)=v_{k}(0)+dv_{k}(0)$ remains normalized. We use the SDE from Theorem 3 to compute that up to terms of order $\mathrm{d}t^{3/2}$ or higher we have (omitting “(0)” in the notation)
@@
The argument may be iterated and thus the claim holds for all $t>0$.
@@ Appendix D Time reversal: Existence and Uniqueness
The time-reversal of the (Dyson-BM) in the sense of Ref. [20] is given in Eq. (4). Since the drift coefficient in the (Dyson-BM) is not locally Lipschitz continuous, the existence and uniqueness of strong solutions to (Dyson-BM) and Eq. (4), and the applicability of Ref. [20] are not obvious. While the existence and uniqueness of a strong solution to the (Dyson-BM) is well established (see, e.g., Lemma 4.3.3 in Ref. [47]), we here sketch how to ensure the other two points.
To ensure existence and uniqueness of a strong solution to Eq. (4), repeat the arguments in Lemma 4.3.3 in Ref. [47], where the divergent $x^{-1}$-term in the drift is replaced by the locally Lipschitz continuous approximation $\phi(x)=x^{-1}$ for $|x|\geq R^{-1}$ and $\phi(x)=xR^{2}$ otherwise. For any $R>0$, the desired statements follow from the local Lipschitz continuity of the drift, and details on the limit $R\to 0$ can be found in Ref. [47]. The only difference to the forward motion is the additional drift term containing the score function. However, since the dynamics is almost surely contained in the interior of the Weyl chamber [48], the propagator in the score-contribution is, as usual, dominated by white noise as $\mathrm{d}t\to 0$. Therefore, this term will not cause complications as $R\to 0$ and the arguments from Ref. [47] imply uniqueness and existence of a strong solution to Eq. (4).
Using the same regularization of the divergent term, for any $R>0$ the statements of Ref. [20] are directly applicable. Since this regularization is piecewise continuous, we can take the limit $R\to 0$ under time reversal. Since we just established the existence of the solution to Eq. (4), the limit $R\to 0$ converges. Furthermore, the arguments are expected to generalize (regularize $\phi(x)=x^{-2}$) to the eigenvector equations and their time reversal.
@@ Appendix E Time rescaling
We rescale time for (Dyson-BM). Let $T(s)$ be a continuous, differentiable rescaling of time, monotonically increasing, with $T(0)=0$ (for convenience).
Let $\gamma_{k}(t):=\lambda_{k}(T(t))$. With this definition, the Ito SDE Equation˜Dyson-BM corresponds to the Ito integral
@@
$=$ $\int_{0}^{t}\left(\alpha\sum_{i\neq k}\frac{1}{\lambda_{k}(T(s))-\lambda_{i}(T(s))}-\beta\lambda_{k}(T(s))\right)T^{\prime}(s)\mathrm{d}s+\int_{0}^{t}\sqrt{2\alpha T^{\prime}(s)}\mathrm{d}B_{k}(s)$ (using Thm. 8.5.7 in [49])
@@
Dyson’s conjecture says that the $\lambda(\frac{1}{N},\frac{1}{2})$-SDE converges to global equilibrium in time $\Theta(1)$ (see [27]). Running $\lambda(\frac{1}{N},\frac{1}{2})$ until time 1 is the same as running the $\gamma(\frac{N}{2})$-SDE until time $T(1)=\frac{1}{N}$.
@@ Appendix F Stepsize Controller
Dyson Brownian motion almost surely never crosses the singularities. Hence, conditioning on non-crossing corresponds to conditioning on an event of probability one, which does not change the dynamics. Given the noise, we can thus calculate the maximal step size, beyond which we would cross the singularity. This is a very useful upper bound, which we employ in practice to get the numerical scheme working. It has two effects: (1) close to the boundary of the Weyl Chamber, it avoids numerically stepping over the singularities and (2) far from the boundary, it allows for larger step size, increasing efficiency.
@@ F.1 Forward in Time
3:   $\delta t_{k}\leftarrow$ maximal step size based on $\lambda_{k+1}-\lambda_{k}$ and samples $u_{k}$, $u_{k+1}$ as described in Section˜F.1
@@
Figure 5: Forward step size controller which exploits that non-crossing of paths happens with probability 11. The exact calculations are carried out in Section˜F.1.
@@ F.2 Backward in Time
Backwards in time, we carry out the analogous computation for the more involved backwards SDE in Equation˜4. In essence, this boils again down to solving a quadratic equation and considering all edge cases.
@@ Appendix G Shooting mechanism
During inference, we require access to the learned score. In the forward diffusion, solely numerical errors due to the time discretisation may lead to crossings of singularities. On the one hand, when going backwards in time, due to inconsistencies in the learned score, the repulsion in Equation˜4 might be too weak and the sample path may leave the Weyl Chamber for any sensible step size. Since the score is not defined outside the Weyl Chamber, this is problematic. On the other hand, the step size obtained by conditioning on not leaving the Weyl Chamber in this ill-trained point would be so small that the numerical solver would get stuck. To overcome this, we incorporate a shooting mechanism: If the repulsion force is too weak to prevent crossing of the singularity, resulting in a microscopic step size upon conditioning to remain in the Weyl Chamber, we repel with the invariant-state drift, i.e., we replace the learned score with the score in the invariant state, as we explain below. This mechanism ensures that we stay in the Weyl Chamber, while minimizing its impact on the distribution. With well-tuned parameters, the shooting gets rarely triggered (less than $0.5\%$ of steps) but is essential, as already a single event would cause getting stuck (upon conditioning) or leaving the Weyl Chamber.
@@
In the forward dynamics, if a certain step size would lead to the probability-0 event of leaving the Weyl-Chamber, we know that the source of the error is the finite-time-step discrete approximation, so that decreasing the step size will always provide a fix. In the backward dynamics, however, this is not guaranteed: As discussed, a possible reason for this probability-0 event in the backwards dynamics is that the score $s_{\theta}$ might be not perfectly learned: $s_{\theta}(\lambda,t)\neq s(\lambda,t)$ for some $\lambda,t$. To avoid this probability-0 event, we use the following “shooting mechanism”: If we were to leave the Weyl-Chamber, we replace the learned score with the analytically known score in the invariant state, eliminating the use of the neural network at that point and leading effectively to a repulsion with the negative forward drift. This mechanism is not expected to change the dynamics in any unfavorable way since it is only applied very rarely, and only in cases where the actual learned dynamics is a much worse approximation (since it would give rise to measure zero events).
G.1 Comparison to direct sampling with eigendecompositions
One may wonder why we do not directly sample from the matrix-valued process (1) at time $t$, perform an eigendecomposition to obtain $\lambda(t)$, and then use the increment $\lambda(t+\mathrm{d}t)$ to learn the score via Eq. 5. While this would render learning simulation-free—relegating the Dyson SDE to the loss derivation and backward dynamics—we avoid this approach due to computational efficiency and precision. This strategy is approximately $100$ times slower for graphs of size $n=10$, as accurate eigendecomposition is computationally expensive; performing it at every step becomes prohibitively costly. Moreover, because we train on increments along the simulated trajectory, a single simulation run allows us to learn from the entire path up to time $t$. Our SDE implementation renders the training runtime on the order of a simulation-free diffusion model.
Appendix H Complexity dependence on graph size $n$ of numerical step for (Dyson-BM) and (Eigenvector-SDE)
When simulating the (Dyson-BM) numerically, for a fixed eigenvalue $\lambda_{k}$, we need to calculate its distance to any $\lambda_{\ell}$ for $\ell\in[n]\setminus\{k\}$ in order to obtain the repulsion force entering the drift of the (Dyson-BM). Since we need to do this for each component, we have for a fixed time step a complexity of $\mathcal{O}\left(n^{2}\right)$.
When simulating (Eigenvector-SDE) numerically, for one time step, we update the entire set of $n$ eigenvectors together. Each of the n(n1)/2n(n-1)/2 entries of the Lie algebra can be computed in time $\Theta\left(1\right)$ (note that $E_{(l,k)}$ is a basis element of the Lie algebra). This update requires thus $\mathcal{O}\left(n^{2}\right)$. Accounting for the projection from the Lie algebra in the Lie group, this yields a complexity of $\mathcal{O}\left(n^{3}\right)$ for an eigenvector update.
Since (Dyson-BM) decouples from (Eigenvector-SDE), we thus have a complexity for an update step of $\mathcal{O}\left(n^{2}\right)$ if one is interested in only the spectrum, and a complexity of $\mathcal{O}\left(n^{3}\right)$ if one is interested in the spectrum and eigenvectors. Note that this modularity allows the practitioner to first learn the spectrum in isolation and do any hyperparameter tuning steps on the (smaller) spectral model, and then use this already tuned model to train and optimize for the remaining eigenvector dynamics.
In comparison, an OU-based approach on the entire graph with a GNN-based algorithm requires for a fixed time point an update of the SDE in $n^{2}$ entries in time $\mathcal{O}\left(n^{2}\right)$ followed by $C$ message passing steps (a GNN is trained using message passing, where $C$ is a constant defining how many hops a message is being passed), with each message passing taking time $\mathcal{O}\left(n^{3}\right)$, leading to an overall complexity of $\mathcal{O}\left(n^{3}\right)$ per time step, assuming that $C$ is constant.
Appendix I Empirical Considerations
While we give a thorough analysis of the computational complexity of each update step, the number of time steps is determined by the adaptive step size algorithm. For practical reasons, our method can be configured with a minimal (and maximal) step size. We illustrate the distribution of the number of time steps on the Brain dataset in Figure˜6, where it can be observed that the number of time steps taken concentrates very well. Note that this number of time steps perfectly suffices even for a dataset as large as Brain, which contains $15,000$ graphs. Note further that while these are the number of steps the numerical integrator takes forward, learning has to happen only on much fewer steps, through the described pre-defined schedule. In the case of Brain, learning happens only on $651$ time points. The numerical generation of $105,000$ paths with step number depicted in Figure˜6 took in total $5$ seconds. For memory considerations, note that we require for training only a single A100 GPU ($80$ GB Video Random Access Memory).
We further note that the skip event (see Algorithm˜1) is indeed a rare event: In the Brain setting, the empirical probability of a step being a skip is $10^{-5}$. Similarly, shootings in the analogous backwards runs happen rarely. Note that skips and shoots are needed for correctness and stability, but no negative impact on runtime due to their rarity.
Figure 6: Concentration of time steps of the numerical integrator: Within 55 second, we generated $105,000$ forward paths using our numerical scheme, where the total number of steps per path are distributed as depicted in this histogram. As can be seen, the distribution concentrates well.
Appendix J Inference